# Hypergraphs as Weighted Directed Self-Looped Graphs: Spectral Properties, Clustering, Cheeger Inequality

### v1

@@ B.3 Global Partitioning Metrics
TP(True $\begin{split}&\textit{TP(True Positive)}=|\mathcal{A}_{m}\cap\mathcal{A}_{c}|\\ &\textit{FP(False Positive)}=|\mathcal{A}_{m}\setminus\mathcal{A}_{c}|\\ &\textit{FN(False Negative)}=|\mathcal{A}_{c}\setminus\mathcal{A}_{m}|\\ &\textit{precision}=\frac{TP}{TP+FP}\\ &\textit{recall}=\frac{TP}{TP+FN}\\ &\textit{F1}=\frac{2\times\textit{precision}\times\textit{recall}}{\textit{precision}+\textit{recall}}\end{split}$ (56)
@@
Specifically, for 2-way NCut, we have another equivalent Definition 10. Assume the actual labeled classes are $\mathcal{V}_{1},\mathcal{V}_{2},...,\mathcal{V}_{k}$, where $\mathcal{V}_{i}$ is all the vertices in class $i$. We first compute the F1 scores between $\mathcal{S}_{i}$ and $\mathcal{V}_{j}$ for $i,j\in\{1,2,...,k\}$, then greedily match the resulted sets with the actual labeled classes [29]. For example, if we have three classes with the F1 matrix, where $\mathbf{F}_{i-1,j-1}$ is the F1 score of $\mathcal{S}_{i}$ and $\mathcal{V}_{j}$
@@
Then, we first match $\mathcal{S}_{1}$ with $\mathcal{V}_{2}$ because they have highest F1 score of 0.9. Then we match $\mathcal{S}_{2}$ with $\mathcal{V}_{1}$ because they have the second largest F1 score of 0.8. Then we try to match $\mathcal{S}_{3}$ with $\mathcal{V}_{2}$ because they have a third largest F1 score of 0.7, but $V_{2}$ has already been matched with $\mathcal{S}_{3}$, so we cannot match it again. Then, we try to match $\mathcal{S}_{3}$ with $\mathcal{V}_{3}$ and this time none of $\mathcal{S}_{3}$ and $\mathcal{V}_{2}$ has been matched. As all the sets have been matched, the greedy-match algorithm ends.
@@
B.4 From 2-way Clustering to k-way Clustering
We have two strategies from 2-way clustering to k-way clustering. For the datasets that each actual labeled class has a similar number of instances, we apply 2-way clustering on the largest cluster every time to split it into two clusters. For k-way clustering, this process calls the 2-way clustering k-1 times.
@@ B.4 From 2-way Clustering to k-way Clustering
B.5 Baselines
CLIQUE++. For each hyperedge $e$, each $u,v\in e$ with $u\neq v$, we add an edge $uv$ of weight $w(e)$. Then, we compute the adjacency matrix $A$ and degree matrix $D$. We calculate the graph Laplacian matrix by $L=D-A$. Finally, we do eigen-decomposition for the random walk Laplacian $L_{RW}=D^{-1}L$, whose eigenvalues are associated with the graph NCut value [22]. We calculate the eigenvector associated with the second smallest eigenvalue for global partitioning: we put the non-negative entries as one cluster and the negative entries as another.
@@ B.5 Baselines
DiffEq [41]. We directly use the official code [^3] https://github.com/atsushi-miyauchi/Hypergraph_clustering_based_on_PageRank of this algorithm. This method sweeps over the sweep sets obtained by differential equations for local clustering. For global partitioning, it simply calls local clustering for every vertex and returns the best in terms of conductance. Originally, this algorithm could only take one starting vertex for local clustering. We modified the code to add one additional vertex that connects the 5 starting vertices in each observation. Then regard the newly added vertex as the starting vertex, so that it can obtain the local cluster for the given 5 starting vertices.
node2vec [21]. node2vec embeds a graph using random walks. Since we have the random walk matrix $P$ on the hypergraph, we construct a directed weighted graph by $P$ as the input of node2vec. However, given that there are too many non-zero entries in $P$, node2vec is extremely slow. Therefore, after we get $P$, we only keep the entries of $P$ that are larger than a small threshold. This threshold is tuned so that the (1) execution time is acceptable; (2) the NCut value and F1 value of the result are both near convergence. We did not modify other default hyperparameters in node2vec. After we obtain the vertex embedding, we call KMeans from scikit-learn to directly obtain k clusters for k-way clustering.
event2vec/hyperedge2vec [16]. We directly use the official code [^4] https://github.com/guoji-fu/Event2vec of this algorithm. event2vec was originally designed for heterogeneous graphs [50, 45, 51]. For example, in a citation network that has publications and authors, each publication is an event. We notice that here an event is equivalent to a hyperedge. We first convert the hypergraph into a STAR graph, then we regard each added vertex as an event and call event2vec to obtain vertex embeddings. We tune the training epochs near default so that (1) execution time is acceptable; (2) the NCut value and F1 value of the result are both near convergence. We did not modify other default hyperparameters in event2vec. After we obtain the vertex embedding, we call KMeans from scikit-learn to directly obtain k clusters for k-way clustering.
B.6 Supplementary Experiment Data
B.6.1 Standard Deviations of Nondeterministic Algorithms on Global Partitioning
@@
B.6.2 Time Comparison on Global Partitioning
@@
HyperClus-GBest of Python $\frac{\text{HyperClus-G}}{\text{Best of Python Baselines}}$ ratio 24.34% 37.83% 99.27% 64.69% 49.96% 94.59% 35.67% 123.7% 44.86%

### v2

@@ C.3 Global Partitioning Metrics
$\begin{split}&\textit{TP(True Positive)}=|\mathcal{A}_{m}\cap\mathcal{A}_{c}|\\ &\textit{FP(False Positive)}=|\mathcal{A}_{m}\setminus\mathcal{A}_{c}|\\ &\textit{FN(False Negative)}=|\mathcal{A}_{c}\setminus\mathcal{A}_{m}|\\ &\textit{precision}=\frac{TP}{TP+FP}\\ &\textit{recall}=\frac{TP}{TP+FN}\\ &\textit{F1}=\frac{2\times\textit{precision}\times\textit{recall}}{\textit{precision}+\textit{recall}}\end{split}$ (57)
@@
Specifically, for 2-way NCut, we have another equivalent Definition 16. Assume the actual labeled classes are $\mathcal{V}_{1},\mathcal{V}_{2},...,\mathcal{V}_{k}$, where $\mathcal{V}_{i}$ is all the vertices in class $i$. We first compute the F1 scores between $\mathcal{S}_{i}$ and $\mathcal{V}_{j}$ for $i,j\in\{1,2,...,k\}$, then greedily match the resulted sets with the actual labeled classes (Kollias et al., 2012). For example, if we have three classes with the F1 matrix, where $\mathbf{F}_{i-1,j-1}$ is the F1 score of $\mathcal{S}_{i}$ and $\mathcal{V}_{j}$
@@
Then, we first match $\mathcal{S}_{1}$ with $\mathcal{V}_{2}$ because they have the highest F1 score of 0.9. Then we match $\mathcal{S}_{2}$ with $\mathcal{V}_{1}$ because they have the second largest F1 score of 0.8. Then we try to match $\mathcal{S}_{3}$ with $\mathcal{V}_{2}$ because they have a third largest F1 score of 0.7, but $V_{2}$ has already been matched with $\mathcal{S}_{3}$, so we cannot match it again. Then, we try to match $\mathcal{S}_{3}$ with $\mathcal{V}_{3}$ and this time none of $\mathcal{S}_{3}$ and $\mathcal{V}_{2}$ has been matched. As all the sets have been matched, the greedy-match algorithm ends.
@@
C.4 From 2-way Clustering to $k$-way Clustering
We have two strategies from 2-way clustering to $k$-way clustering. For the datasets that each actual labeled class has a similar number of instances, we apply 2-way clustering on the largest cluster every time to split it into two clusters. For $k$-way clustering, this process calls the 2-way clustering kk-1 times.
@@ C.4 From 2-way Clustering to $k$-way Clustering
Another promising future direction is to develop a direct spectral analysis for the k-way case. This, however, involves non-trivial technical challenges and lies beyond the scope of the present work. We note that such an extension should also draw inspiration from analogous results in graphs, such as the k-way NCut formulation.
Definition 25 (k-way Normalized Cut).
Let $H=(V,E,\omega,\gamma)$ be an EDVW hypergraph with stationary distribution $\phi$ and transition matrix $P$. For any $k$-way partition $\{S_{1},\dots,S_{k}\}$ of $V$ into non-empty disjoint subsets, the $k$-way normalized cut (NCut) is defined as
$NCut(S_{1},\dots,S_{k})\;=\;\sum_{i=1}^{k}\frac{|\partial S_{i}|}{vol(S_{i})}\;=\;\sum_{i=1}^{k}\frac{\displaystyle\sum_{u\in S_{i},\,v\in\overline{S_{i}}}\phi(u)P_{uv}}{\displaystyle\sum_{u\in S_{i}}\phi(u)},$ (61)
where $vol(S_{i})=\sum_{u\in S_{i}}\phi(u)$ is the stationary volume of cluster $S_{i}$, and $|\partial S_{i}|=\sum_{u\in S_{i},\,v\in\overline{S_{i}}}\phi(u)P_{uv}$ denotes the boundary volume between $S_{i}$ and its complement.
C.5 Baselines
CLIQUE++. For each hyperedge $e$, each $u,v\in e$ with $u\neq v$, we add an edge $uv$ of weight $w(e)$. Then, we compute the adjacency matrix $A$ and degree matrix $D$. We calculate the graph Laplacian matrix by L=DAL=D-A. Finally, we do eigen-decomposition for the random walk Laplacian $L_{RW}=D^{-1}L$, whose eigenvalues are associated with the graph NCut value (Hamilton, 2020). We calculate the eigenvector associated with the second smallest eigenvalue for global partitioning: we put the non-negative entries as one cluster and the negative entries as another.
@@ C.5 Baselines
DiffEq (Takai et al., 2020). We directly use the official code [^4] https://github.com/atsushi-miyauchi/Hypergraph_clustering_based_on_PageRank of this algorithm. This method sweeps over the sweep sets obtained by differential equations for local clustering. For global partitioning, it simply calls local clustering for every vertex and returns the best in terms of conductance. Originally, this algorithm could only take one starting vertex for local clustering. We modified the code to add one additional vertex that connects the 5 starting vertices in each observation. Then regard the newly added vertex as the starting vertex, so that it can obtain the local cluster for the given 5 starting vertices.
node2vec (Grover & Leskovec, 2016). node2vec embeds a graph using random walks. Since we have the random walk matrix $P$ on the hypergraph, we construct a directed weighted graph by $P$ as the input of node2vec. However, given that there are too many non-zero entries in $P$, node2vec is extremely slow. Therefore, after we get $P$, we only keep the entries of $P$ that are larger than a small threshold. This threshold is tuned so that the (1) execution time is acceptable; (2) the NCut value and F1 value of the result are both near convergence. We did not modify other default hyperparameters in node2vec. After we obtain the vertex embedding, we call KMeans from scikit-learn to directly obtain k clusters for $k$-way clustering.
event2vec/hyperedge2vec (Fu et al., 2019). We directly use the official code [^5] https://github.com/guoji-fu/Event2vec of this algorithm. event2vec was originally designed for heterogeneous graphs. For example, in a citation network that has publications and authors, each publication is an event. We notice that here an event is equivalent to a hyperedge. We first convert the hypergraph into a STAR graph, then we regard each added vertex as an event and call event2vec to obtain vertex embeddings. We tune the training epochs near default so that (1) execution time is acceptable; (2) the NCut value and F1 value of the result are both near convergence. We did not modify other default hyperparameters in event2vec. After we obtain the vertex embedding, we call KMeans from scikit-learn to directly obtain k clusters for $k$-way clustering.
C.6 Supplementary Experiment Data
C.6.1 Standard Deviations of Nondeterministic Algorithms on Global Partitioning
@@
C.6.2 Time Comparison on Global Partitioning
@@
$\frac{\text{HyperClus-G}}{\text{Best of Python Baselines}}$ ratio 24.34% 37.83% 99.27% 64.69% 49.96% 94.59% 35.67% 123.7% 44.86%